OpenStudy (anonymous):

What is the vertex for the graph of y = -2x^2 + 8x - 13? Answer (2, 11) (2, -5) (-2, -37) (-2, -21)

OpenStudy (anonymous):

Lets graph this! (2,-5)

OpenStudy (anonymous):

vertex is x=-b/2a x = -8 /2(-2) x = -8 / -4 x = 2 y = -2(2)^2 + 8(2) - 13 y = -2(4) + (16) - 13 y = -8 + 16 - 13 y = -5 vertex point: (2, -5) This is a good graphing calculator and can help on these types of questions: https://www.desmos.com/calculator

OpenStudy (anonymous):

Its much easier to do just graph it like @Mrbonez321 though:) So as long as you have a graphing calculator you can avoid the formulas at least for a while:)